OpenStudy (anonymous):

one more and i promise ill give up for the night (; Write the equation of a circle with center at (1, 2) and a diameter of 20. A. x2 + (y – 2)2 = 100 B. (x - 1)2 - (y – 2)2 = 100 C. (x - 1)2 + (y – 2)2 = 100 D. (x - 1)2 + y2 = 100

OpenStudy (anonymous):

center is (1, 2) so we're looking for (x - 1)^2 + (y - 2)^ = r^2 radius is the half the diameter r = 10 ====> r^2 = 100 so (x - 1)^2 + (y - 2)^ = 100 looks like C

OpenStudy (primeralph):

@robz8 on point.

OpenStudy (anonymous):

damn, i thought it was B. glad i asked. thanks for all your help man, i really appreciate it. sometimes ya gonna lend a brain, and im glad you were there for me. (:

OpenStudy (primeralph):

@chicagoblkhaks you don't have to give up......keep them coming.

OpenStudy (anonymous):

its my bed time. ): guys. this little lady needs her beauty sleep. haha.

OpenStudy (anonymous):

... bruins...

OpenStudy (primeralph):

Okay then. Good night.

OpenStudy (anonymous):

hey now, im actually a wild fan. but my boys gave that up. ill root for your bruins though, id like to see them win to be honest!!

OpenStudy (anonymous):

that's funny... yet your name says blackhawks... :P